Use the substitution method to solve the linear system given. Express
the answer as an ordered pair of the form (x,y)
x + y = 5
2x + y = 6

Answer by checkley77(12844)

x + y = 5 or y=5-x
2x + y = 6
Now replace y in the second equation with (5-x) & solve for x.
2x+(5-x)=6
2x+5-x=6
x=6-5
x=1 ans.
y=5-1
y=4 ans.
Proof:
2*1+4=6
2+4=6
6=6